Consider a competitive market served by many domestic and foreign firms. The domestic demand for these firms’ product is Qd = 1300 - 1.5P. The supply function of the domestic firms is QSD = 150 + 0.5P, while that of the foreign firms is QSF = 200.

Instructions: Enter your responses for equilibrium price rounded to the nearest penny (two decimal places). Enter your responses for equilibrium quantity rounded to one decimal place.

a. Determine the equilibrium price and quantity under free trade.

Equilibrium price: $

Equilibrium quantity: units


b. Determine the equilibrium price and quantity when foreign firms are constrained by a 100-unit quota.

Equilibrium price: $

Equilibrium quantity: units


c. Are domestic consumers better or worse off as a result of the quota?

  • Better off

  • Neither better nor worse off

  • Worse off



d. Are domestic producers better or worse off as a result of the quota?

  • Better off

  • Neither better nor worse off

  • Worse off
Answer #1

a) QD = 1300-1.5P

QS = 150+0.5P+200 = 350+0.5P

QD=QS

1300-1.5P = 350+0.5P

950 = 2P

P = 950/2 = 475

Q = 1300-1.5(475) = 587.5

b) QD=1300-1.5P

QS= 150+0.5P+100 = 250+0.5P

QD=QS

1300-1.5P = 250+0.5P

1050 = 2P

P = 1050/2 = 525

Q = 1300-1.5(525) = 512.5

c) Consumers are worse off

d) Producers are better off